pq=3x+14. qr=5x−44 Find x


1
Expert's answer
2021-01-10T17:09:40-0500

Assuming Q is the midpoint of line PQ and QR, we can say:

3x + 14 = 5x - 44

3x - 5x = -44 - 14

-2x = -58

x = (58/2) = 29.